OpenStudy (anonymous):

What is the coordinate of the center of the circle given by the equation (x-3)^2+y^2=15 A. (3,0) B. (-3,0) C. (-3,15) D. (0,3)

OpenStudy (anonymous):

\[ (x-3)^2+(y-0)^2=15\]compare to \[ (x-h)^2+(y-k)^2=r^2\]

OpenStudy (jdoe0001):

\(\bf (x-3)^2+y^2=15\implies (x-{\color{blue}{ 3}})^2+(y-{\color{blue}{ 0}})^2=15\qquad {\color{blue}{ center}}\)

OpenStudy (anonymous):

(h,k) is the center :)

OpenStudy (anonymous):

Thank yall for the help

OpenStudy (jdoe0001):

yw